A die has been loaded sop that the probability of rolling an even number is 2/9 and the probability of rolling an odd is 1/9. What is the probability of.
(i) rolling a 6, given that an even number is rolled
(ii) rolling a 3, given that an odd number is rolled.
1
Expert's answer
2020-07-02T19:31:40-0400

The probability of.

(i) rolling a 6, given that an even number is rolled

=(2/9)/(6/9)=2/6=1/3

(ii) rolling a 3, given that an odd number is rolled.

=(1/9)/(3/9)=1/3